OpenStudy (anonymous):

Solve the following systems of equations.. Just when I thought i was done lol another assignement SHOOT ME Lol... x^2+y^2=26 y-x=-6

OpenStudy (anonymous):

Please help I really need to get this done!! PLEASE

OpenStudy (anonymous):

Nm dont worry about this one... figured it out.. i might ahve a new question shortley

OpenStudy (anonymous):

x^2+y^2=21 x^2-y=9

OpenStudy (anonymous):

square the second equation, x^2+y^2+2xy=36. 26+2xy=36 xy=5 x=5/y y-5/y+6=0 y^2+6y-5=0 .

OpenStudy (anonymous):

I woudl be really thankful if you can do that one i ust sent x^2+y^2=21 x^2-y=9

OpenStudy (anonymous):

9+y+y^2=21 y^2+y-12=0; y=-4,y=3 then find x.

OpenStudy (anonymous):

how?? i suck at these

OpenStudy (anonymous):

can you tell me x plz?

OpenStudy (anonymous):

if y=3,x=sqrt(12) if y=-4,x=sqrt(5).

OpenStudy (anonymous):

All i did was i substitued x^2 from the secnd equation onto the first.

OpenStudy (anonymous):

it says there are still more solutions??

OpenStudy (anonymous):

can you help or shuld i repost?